# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ 2 Literature Review
    Which AI/ML techniques are commonly employed in supply chain risk assessment, prediction, and mitigation?
@@
3 Methodology
3.1 Study Design
The research utilized a systematic approach to conduct a comprehensive literature review coupled with a bibliometric analysis. This methodology aimed to scrutinize the intersection of AI/ML techniques and SCRA, providing insights into research trends, methodologies, and emerging themes.
The SLR followed a robust methodology, adhering to the guidelines proposed by (Denyer and Tranfield,, 2009). This approach aimed to identify and analyze research articles focusing on integrating AI/ML techniques within SCRA.
@@ 3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS). We used the ”Publish or Perish 8” tool for the Google Scholar search to filter articles published between 2015 and 2023. The initial search yielded 1000 articles. In the WoS search, we employed the following search strings to retrieve relevant articles:
(ALL=((”supply chain”) AND (”risk*” OR ”credit risk*” OR ”risk assessment*” OR ”risk prediction” OR ”disruption*” OR ”disease outbreak*” OR ”post COVID resilienc*”) AND (”artificial intelligence” OR ”machine learning” OR ”neural network*” OR ”deep learning” OR ”reinforcement learning” OR ”SVM” OR ”support vector machine” OR ”boosting” OR ”ensemble” OR ”bayesian network model” OR ”random forest” OR ”LSTM” OR ”long short-term memory”))) NOT (DT==(”PROCEEDINGS PAPER” OR ”BOOK CHAPTER” OR ”EDITORIAL MATERIAL” OR ”LETTER”))
The WoS search initially retrieved 434 articles.
@@ 3.3 Article Selection Process
We followed a systematic approach to select the most relevant articles for our analysis. The inclusion and exclusion criteria were applied sequentially to ensure the selection of high-quality research. The following steps were taken:
  • 1.
    Removal of Duplicates: We removed duplicate articles identified between the two databases, resulting in 360 duplicates.
  • 2.
    Filtering by Reliable Publishers: We only considered articles published by reliable and reputable publishers, excluding 398 articles.
  • 3.
    Non-English Exclusion: We excluded articles not written in English, resulting in removing 3 articles.
  • 4.
    Filtering by Publication Type: We excluded articles classified as Q3 or Q4, proceedings papers, book chapters, editorials, and letters, resulting in the removal of 244 articles.
  • 5.
    Title, Abstract, and Conclusion Review: We reviewed the remaining articles based on their title, abstract, methodology, and conclusion to ensure their relevance to our research questions.
This step removed 386 articles, leaving us with a final set of 48 research articles.
@@ 3.4 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis.
To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 434 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections.
Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2023, and it is possible that relevant articles published before or after this timeframe were not included.
@@ 4 Bibliometric Analysis
In our quest to comprehensively analyze the landscape of supply chain risk assessment using machine learning, we employed an array of tools, including MS Excel and VOSviewer software. This bibliometric analysis comprised several key components, each contributing to a deeper understanding of the field.
@@ 4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2014-2023) of the reviewed literature.
Figure 2 offers a dynamic portrayal of the publication trend over the years, presenting an invaluable temporal dimension to our bibliometric analysis. A meticulous examination of data showed how the research landscape has evolved from 2014 to 2023. This longitudinal perspective equips us with the knowledge of emerging trends, shifts in focus, and the growth trajectory of the field.
@@ 4.2 Geographical Distribution of Research
In Figure 3, we present a geographical distribution of research related to SCRA using machine learning. The figure offers a clear visualization of countries that have actively contributed to the literature in this field. The deeper coloration signifies a higher volume of research publications from these regions. This representation enriches our understanding of the global landscape of SCRA, highlighting regions with significant scholarly contributions.
@@ 4.3 Analysis of Authors
Table 1: Top 10 authors according to the published papers
Authors Published Paper Count
Gupta S 10
Kumar A 8
Liu YK 7
Brintrup A 5
Liu C 5
Modgil S 5
Yang M 5
Bouzembrak Y 4
Choi TM 4
Chung SH 4
Table  1 presents a ranking of the top 10 authors in the field of risk prediction and analysis based on the number of research papers attributed to each. ”Gupta S” leads the list with 10 papers, highlighting a substantial body of work in this domain. This table serves as a concise reference for identifying influential authors who have made significant contributions to the field of risk analysis and provides valuable insights into their scholarly output, aiding researchers and professionals in exploring their work.
@@ 4.4 Publisher Contributions
Figure 4 provides a compelling visual representation of the leading publishers in the field of SCRA, with a specific emphasis on applications of AI and ML. This chart clearly depicts the number of research papers associated with each publisher, enabling researchers and professionals to discern the most prolific contributors to this dynamic domain. It serves as a valuable resource for identifying key publishers that disseminate influential research at the intersection of AI/ML and SCRA, granting a panoramic view of the publication landscape in this evolving field.
@@ 4.5 Journal Contributions
This bar chart, represented in Figure 5, showcases the top 11 journals significantly contributing to the field of SCRA with a specific focus on AI and ML applications. This visual representation helps researchers and professionals identify the key journals for accessing relevant and impactful research in AI/ML and SCRA. It concisely overviews this dynamic field’s most prolific journal sources.
4.6 Co-citation Analysis
In pursuing a comprehensive understanding of the scholarly landscape in AI-based SCRA, we present three co-citation analyses derived from bibliometric data of the authors, cited references, and publication sources using VOSviewer. Each visualization offers unique insights into collaboration, thematic prominence, and citation patterns within this dynamic field.
@@ 4.6 Co-citation Analysis
(c) Co-citation network graph of publication sources
@@
Figure 5(a) illustrates connections between authors, highlighting 55 out of 18,858 authors surpassing a minimum citation threshold of 20. These authors formed three distinct clusters based on their co-citation patterns: Red (22 items), Green (19 items), and Blue (14 items). The clusters signify groups of authors who are frequently cited together, suggesting shared research themes or collaboration networks. Node size signifies citation counts, and connections depict co-citation relationships. This visualization unveils collaborative networks and shared contributions among influential authors.
The co-citation analysis of cited references involved a minimum citation threshold of 10, identifying 58 cited references meeting this criterion (Figure 5(b)). These references formed five distinct clusters, reflecting thematic relationships among the cited works. The clusters shed light on the interconnectedness of references, providing insights into key themes or influential works within the field. Five formed clusters and individual nodes within each analysis were carefully examined for meaningful patterns. Each cluster represents distinct research themes or methodologies, while individual nodes signify specific authors, countries, or references that play pivotal roles in the overall co-citation network.
A minimum of 50 citations per source was set to analyze publication sources, identifying 70 sources meeting the criteria (Figure 5(c)). These sources were grouped into three clusters based on their co-citation patterns. These clusters offer insights into the thematic literature concentrations within your field of study. The three formed clusters were examined to discern the major themes and sub-fields represented by the co-cited sources. The significance of central sources was considered within each cluster, as they represented foundational works or widely acknowledged references in the field.
@@ 4.7 Co-authorship Analysis
Co-authorship analysis is a method to explore and understand collaborations among authors or researchers based on their joint publication activities. This analysis highlights the collaborative essence inherent in this research domain, revealing influential partnerships among authors and the noteworthy contributions of specific countries.
(a) Co-authorship network analysis of 63 authors.
(b) Co-authorship network analysis of 47 countries.
Figure 7: Co-authorship network graphs visualization representing (a) authors and (b) countries
Figure 6(a) unveils the co-authorship network among authors within this field. By setting a threshold of at least 3 documents per author, we identified 63 authors meeting this criterion. The resulting visualization, created using VOSViewer, showcased 6 distinct colored clusters. For instance, clusters colored red, green, and deep blue comprised 10, 7, and 5 authors, respectively. The clusters are formed based on similar collaborative behaviors among these authors. The links connecting authors within clusters represent the robustness of their collaboration, indicating frequent joint authorship or shared research endeavors.
Figure 6(b) demonstrates co-authorship analysis at a country level. Using data encompassing document count, citations, and total link strength, a minimum document threshold of 2 per country identified 47 countries meeting this criterion. The visualization prominently emphasizes China’s substantial dominance across document count, citations, and total link strength compared to other countries. This dominance underscores China’s extensive contributions to the realm of Supply Chain Risk Assessment utilizing AI/ML techniques.
@@ 4.8 Co-occurrence Analysis
The results of our co-occurrence analysis using VOSviewer, focusing on the unit of analysis as ’Authors’ and ’Keywords Plus,’ were presented in Figure 8, offering insights into the thematic intricacies of SCRA and ML. The co-occurrence analysis is a powerful tool to reveal latent connections and thematic clusters within the expansive landscape of scholarly literature. VOSviewer facilitated the exploration of how certain keywords converge within the existing body of literature. This visualization enables us to discern the prevalent themes and areas of emphasis within this research space, guiding our understanding of the critical topics that have garnered scholarly attention.
(a) Co-occurrence network analysis of top 44 authors keywords.
(b) Co-occurrence network analysis of top 42 keywords used in the literature related to SCRA-AI.
Figure 8: The Co-occurrence Network visualizes the intricate relationships among the identified keywords. Each node represents a keyword, and the edges denote the frequency of co-occurrence. Clusters of tightly connected nodes signify thematic groupings, offering a comprehensive snapshot of the intellectual structure within the analyzed literature.
In Figure 7(a), the network map illustrates the relationships among these keywords based on their co-occurrence within the authorship data. Each node represents a keyword, and the links between nodes signify the strength of co-occurrence. Larger nodes indicate keywords with higher occurrences, while the proximity of nodes reflects the degree of association. In this analysis, 1483 keywords were initially considered, but after applying a minimum threshold of 4 occurrences, 44 keywords emerged as frequent keywords.
In Figure 7(b), the analysis was conducted by setting a minimum threshold of 10 occurrences for each keyword, identifying 42 keywords that surpassed this criterion from a total pool of 1032 keywords. The 42 keywords meeting the established threshold were subjected to clustering analysis, unveiling intricate co-occurrence patterns. These 4 clusters signify prevalent themes and provide insights into the interconnectivity of concepts within our research domain.
@@ 4.9 Top 10 Affiliations
Table 2: Top 10 Affiliations According to Paper Count
Affiliations Published Paper Count % of 434
INDIAN INSTITUTE OF TECHNOLOGY SYSTEM IIT SYSTEM 13 2.995
HONG KONG POLYTECHNIC UNIVERSITY 11 2.535
NEOMA BUSINESS SCHOOL 10 2.304
CHINESE ACADEMY OF SCIENCES 9 2.074
EMLYON BUSINESS SCHOOL 9 2.074
NATIONAL INSTITUTE OF INDUSTRIAL ENGINEERING NITIE 8 1.843
UNIVERSITY OF CAMBRIDGE 8 1.843
HEBEI UNIVERSITY 7 1.613
INDIAN INSTITUTE OF TECHNOLOGY IIT DELHI 7 1.613
UNIVERSITY OF TECHNOLOGY SYDNEY 7 1.613
OTHERS 345 79.49
Table  2 provides an overview of the top 10 affiliations based on paper count in the field of SCRA. These affiliations have significantly contributed to the literature, collectively accounting for a notable portion of the total papers reviewed (2.995% to 1.613%). The table underscores the involvement of various academic and research institutions in advancing research related to SCRA.

### v3

@@ 3.1 Study Design
    Which AI/ML techniques are commonly employed in SCRA, prediction, and mitigation?
@@
The SLR followed a robust methodology, adhering to the guidelines proposed by Denyer and Tranfield, (2009). This approach aimed to identify and analyze research articles focusing on integrating AI/ML techniques within SCRA.
@@ 3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS). We used the “Publish or Perish 8” (Harzing,, 2016) software for the Google Scholar search to filter articles published between 2015 and 2024. The initial search yielded 1000 articles. In the WoS search, we employed the following search strings to retrieve relevant articles:
(ALL=( “supply chain”) AND (“risk*” OR “credit risk*” OR “risk assessment*” OR “risk prediction” OR “disruption*” OR “disease outbreak*” OR “post COVID resilienc*”) AND (“artificial intelligence” OR “machine learning” OR “neural network*” OR “deep learning” OR “reinforcement learning” OR “SVM” OR “support vector machine” OR “boosting” OR “ensemble” OR “bayesian network model” OR “random forest” OR “LSTM” OR “long short-term memory”))) NOT (DT==(“PROCEEDINGS PAPER” OR “BOOK CHAPTER” OR “EDITORIAL MATERIAL” OR “LETTER”))
Table 1: Inclusion and exclusion criteria and justifications
Criterion Inclusion/Exclusion Justification
Date range 2015–2024 Focuses on recent studies, ensuring relevance to current advancements in AI for SCRA.
Language English only Ensures consistency in analysis and accessibility for an English-speaking academic audience.
Duplicate removal Exclusion Avoids redundancy and ensures each study contributes uniquely to the analysis.
Publisher reliability Exclusion Maintains academic quality by including only studies from reputable, peer-reviewed journals.
Publication type Exclusion of Q3/Q4, proceedings, book chapters, etc. Focuses on high-impact research, prioritizing quality over quantity.
Checking for relevance (title, abstract, and conclusion) Exclusion Ensures alignment with the study’s objectives based on content relevant to SC risk and AI.
@@ 3.3 Article Selection Process
To ensure the selection of high-quality and relevant studies, we followed a systematic, multi-stage screening process guided by the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) framework. The selection process was designed to minimize potential bias and ensure transparency at each stage. The stages are detailed as follows:
  1. 1.
    Removal of Duplicates: To avoid redundant analysis, we first identified and removed duplicate articles between the two databases, excluding 360 duplicate entries. This step is standard in systematic reviews to ensure each study contributes uniquely to the analysis.
  2. 2.
    Filtering by Publisher Reliability: We excluded 398 articles published by sources identified as less reliable, focusing only on those published in reputable, peer-reviewed journals. This criterion helps maintain high academic integrity and ensures that the final selection is based on studies that meet established research quality standards.
  3. 3.
    Non-English Exclusion: Since this review focuses on an English-speaking academic audience, we excluded 3 articles not written in English. This ensures that the included studies are accessible and can be consistently interpreted without translation bias.
  4. 4.
    Filtering by Publication Type: To prioritize peer-reviewed journal articles, we excluded papers classified as Q3 or Q4 and proceedings papers, book chapters, editorials, and letters, removing 244 articles. This criterion ensures that our review is based on high-impact research published in top-quality journals, as lower-ranked journals or other types of publications may not always meet rigorous peer-review standards.
  5. 5.
    Checking for Relevance (Title, Abstract, and Conclusion Screening): The remaining articles were reviewed based on their title, abstract, and conclusions to assess their relevance to the research questions of this study. Articles that did not focus specifically on AI applications in SCRA were excluded. This step eliminated an additional 383 articles, leaving a final sample of 51 articles for in-depth review.
Each criterion was carefully chosen to ensure that only studies of high relevance and quality were included, thus minimizing potential biases. Using dual databases (Google Scholar and WoS) ensured comprehensive coverage, while the sequential exclusion criteria systematically filtered out irrelevant or low-quality studies. To demonstrate further transparency, Table 1 outlines the inclusion and exclusion criteria and their justifications.
3.4 Bias Mitigation
To address concerns about potential bias, we followed a systematic approach that minimizes subjective selection at each stage:
  1. 1.
    By using two widely recognized databases, Google Scholar and WoS, we ensured broad initial coverage of the literature on AI in SCRA, reducing the likelihood of selection bias by capturing a diverse range of studies.
  2. 2.
    Each exclusion criterion was pre-defined and systematically applied, following a step-by-step approach that minimized subjective judgment. This ensures that all articles are assessed based on consistent standards.
  3. 3.
    The PRISMA framework was applied rigorously to eliminate irrelevant studies at multiple stages, further reducing the risk of bias by ensuring that only studies meeting high-quality and relevance standards were retained.
3.5 Final Selection
After applying all criteria, 51 articles were deemed relevant and of high quality for inclusion in this review. The selected articles cover various AI methodologies applied to different SC risk scenarios, providing a comprehensive foundation for our analysis. According to Figure 1, 434 articles were included in the bibliometric analysis to provide a broad understanding of the field, considering various quantitative metrics. The 51 shortlisted articles represent a more focused subset. We selected them based on additional qualitative criteria, such as relevance, methodology, and contribution to the research.
@@ 3.6 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis. To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 434 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections. Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2024, and it is possible that relevant articles published before or after this timeframe were not included.
@@ 4 Bibliometric Analysis
To comprehensively analyze the scholarly landscape of SCRA using AI, we employed a combination of tools, including MS Excel and VOSviewer. This bibliometric analysis encompasses several dimensions, offering a deeper understanding of the field by identifying trends, contributions, and intellectual structures.
@@ 4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2015-2024) of the reviewed literature.
Figure 2 illustrates the annual publication trends from 2015 to 2024. A significant rise in publications is observed post-2020, likely influenced by the COVID-19 pandemic, highlighting the critical need for resilient SC systems. The peak in research output during 2021 and 2022 reflects the heightened importance of AI/ML-driven risk assessment in tackling complex SC challenges. However, the slight decline in subsequent years may indicate a maturing research field, transitioning from exploratory studies to more application-oriented implementations in industry.
@@ 4.2 Geographical Distribution of Research
Figure 3 visualizes the geographical distribution of research contributions. China leads with 164 publications, followed by the United States (74) and Australia (32). This dominance underscores the strong focus of Chinese institutions on AI-driven SCRA, supported by extensive investments in AI research and SC optimization. The global spread of contributions emphasizes the universal relevance of SCRA, with increasing participation from emerging economies.
Table 2: Top 10 authors according to the published papers
{tblr}
column2 = c, hline1-2,12 = -, Authors & Published Paper Count
Gupta S 5
Kumar A 4
Modgil S 4
Bag S 3
Hu XJ 3
Wang GJ 3
Xie C 3
Yao G 3
Zhu Y 3
Antoniou G 2
@@ 4.3 Analysis of Authors
As shown in Table  2, the top 10 authors in the field have significantly influenced the domain of SCRA using AI/ML. “Gupta S” leads with 10 publications, followed by “Kumar A” (8) and “Liu YK” (7). These authors represent thought leaders who have contributed extensively to developing AI-based risk assessment frameworks and methodologies.
@@ 4.4 Publisher Contributions
Figure 4 presents the contributions of various publishers. Elsevier is the most prominent, contributing over 120 articles, followed by Springer Nature and MDPI. These publishers serve as key platforms for disseminating high-quality research, reflecting their alignment with the evolving needs of AI and SCRA scholarship.
@@ 4.5 Journal Contributions
In Figure 5, journal contributions are analyzed, with “Sustainability” being the most active source of publications, followed by “International Journal of Production Research” and “Annals of Operations Research.” These journals provide vital platforms for researchers to share AI-driven SC risk management advancements.
@@
(c) Co-citation network graph of publication sources.
@@
4.6 Co-citation Analysis
Co-citation analysis was conducted to explore the intellectual structure of AI-SCRA research, focusing on authors, references, and publication sources. Figure 6(a) illustrates a network of 55 authors grouped into three clusters, emphasizing the pivotal roles of researchers such as Ivanov and Baryannis, who have significantly shaped the development of AI-driven risk assessment frameworks. Figure 6(b) focuses on cited references, identifying five distinct clusters highlighting recurring themes, including adopting hybrid AI models, developing advanced risk prediction frameworks, and exploring post-COVID SCR strategies. Lastly, Figure 6(c) underscores the centrality of academic journals, with the International Journal of Production Research standing out as a key publication venue, reflecting its substantial impact on the field’s scholarly discourse and research progression.
(a) Co-authorship network analysis of 48 authors.
(b) Co-authorship network analysis of 18 countries.
Figure 7: Co-authorship network graphs visualizing collaborations at (a) author and (b) country levels.
@@ 4.7 Co-authorship Analysis
Co-authorship analysis is a powerful method for understanding collaborations among authors and researchers based on joint publication activities. This analysis highlights the collaborative nature of the research domain, uncovering influential partnerships and the significant contributions of specific countries.
Figure 7(a) illustrates the co-authorship network among authors in this field. Using a threshold of at least one document and a minimum of 55 citations per author, 48 authors out of a total of 304 were identified as significant contributors. The visualization, generated using VOSViewer, reveals four distinct clusters, color-coded as red, green, blue, and yellow. These clusters represent groups of authors with strong collaborative ties. For example, the robust connections surrounding “Modgil” in the red cluster indicate frequent co-authorship and shared research initiatives, reflecting a high degree of collaboration within the cluster.
Figure 7(b) showcases the co-authorship network at the country level. By applying thresholds of at least two documents and nine citations per country, 18 out of 36 countries were included in the analysis. The visualization highlights England and China as leading contributors, dominating in terms of document count, citation impact, and total link strength. This dominance underscores their pivotal roles in advancing AI/ML-driven SCRA research. The results also emphasize strong international collaborations between these countries and other key players, fostering a globally interconnected research ecosystem.
(a) Co-occurrence network analysis of top 26 authors’ keywords.
(b) Co-occurrence network analysis of top 47 keywords used in the literature related to SCRA-AI.
Figure 8: The Co-occurrence Network visualizes the intricate relationships among identified keywords. Each node represents a keyword, and the edges denote the frequency of co-occurrence. Clusters of tightly connected nodes signify thematic groupings, providing a comprehensive snapshot of the intellectual structure within the analyzed literature.
@@ 4.8 Co-occurrence Analysis
The co-occurrence analysis offers valuable insights into the thematic structure of SCRA and AI/ML research by identifying how keywords are interconnected within the scholarly literature. Using VOSviewer, we conducted this analysis by focusing on the units of analysis, such as “Authors’ Keywords” and “Keywords Plus.” This approach reveals latent connections and thematic clusters, enabling a deeper understanding of this domain’s critical topics and prevalent themes.
The network map shown in Figure 8(a) illustrates the relationships among the 26 most frequently occurring author keywords. These keywords were selected from an initial pool of 1,483, applying a minimum threshold of three occurrences. Larger nodes, such as “artificial intelligence” and “machine learning,” indicate higher frequencies. The proximity of the nodes reflects their degree of association, with closely grouped nodes forming thematic clusters. Four clusters emerged, highlighting key themes, including risk management, sustainability, and digital transformation.
Figure 8(b) illustrates the co-occurrence of keywords found in the reviewed literature. The analysis identified 47 frequently occurring terms from a total of 1,032, using a minimum threshold of three occurrences. These keywords were categorized into five distinct clusters, reflecting prevalent themes and connections within the research field. Terms such as “management” and “framework” exhibited the strongest linkages, indicating their crucial roles in connecting different thematic areas. Additionally, other significant clusters highlighted emerging trends, including “supply chain resilience,” “blockchain,” and “sustainability.” This co-occurrence analysis offers a comprehensive overview of the intellectual structure of SCRA research, showcasing both established and emerging themes. The findings emphasize the interconnectivity of concepts and provide guidance for future research directions by identifying critical areas of scholarly focus.
Table 3: Top 10 Affiliations According to Paper Count
{tblr}
column2 = c, column3 = c, hline1-2,13 = -, Affiliations & Published Paper Count % of 51 articles
NEOMA BUSINESS SCHOOL 5 5.319
CHINESE ACADEMY OF SCIENCES 4 4.255
EMLYON BUSINESS SCHOOL 4 4.255
HEFEI UNIVERSITY OF TECHNOLOGY 4 4.255
HONG KONG POLYTECHNIC UNIVERSITY 4 4.255
HUNAN UNIVERSITY 4 4.255
INTERNATIONAL MANAGEMENT INSTITUTE (IMI) KOLKATA 4 4.255
UNIVERSITY OF CAMBRIDGE 4 4.255
UNIVERSITY OF TECHNOLOGY SYDNEY 4 4.255
UNIVERSITY OF HUDDERSFIELD 3 3.191
OTHERS 214 57.45
@@ 4.9 Top 10 Affiliations
Table  3 highlights the top 10 affiliations contributing to SCRA research, ranked by their published paper count. These institutions collectively demonstrate significant academic leadership, with NEOMA Business School leading with 5 publications (5.319% of the total reviewed articles). Several institutions, including the Chinese Academy of Sciences, Emlyon Business School, and Hong Kong Polytechnic University, each contributed 4 papers representing 4.255% of the reviewed literature. These affiliations underscore a global research effort, with prominent contributions spanning Europe, Asia, and Australia. Together, the top affiliations account for a substantial share of foundational studies in SCRA, promoting advancements in the application of AI to mitigate SC risks. Meanwhile, the category labeled “Others” collectively represents the remaining 57.45% of contributions, reflecting a broad and diverse research landscape.